OpenStudy (anonymous):

Can someone check me on this? You have a standard number cube. What is the probability of rolling a number less than 3, and then rolling a prime number? 1/3 1/2 1/36 1/6

OpenStudy (anonymous):

I put 1/6

OpenStudy (anonymous):

probability of rolling a number less thatn 3 = 2/6 probabilyty of rolling a prime number = 4/6 so the probabilty that the problem is asking for is (2/6)*(4/6)

OpenStudy (anonymous):

i got 2/9

OpenStudy (anonymous):

that is no in ur options, i dont know where i made the mistake

OpenStudy (anonymous):

i search for this exercise in the web and they get the same result http://answers.yahoo.com/question/index?qid=20101206014111AACNoCo

OpenStudy (anonymous):

since its a number cube the denominator would be 6

OpenStudy (anonymous):

when it says a number less than three, does it also count the three?

OpenStudy (raden):

the prime numbers = {2,3,5}, there are 3 events so, should the probability of rolling a prime number = 3/6 = 1/2

OpenStudy (raden):

therefore, it is 1/3 * 1/2 = 1/6

OpenStudy (anonymous):

Thank you so much @RadEn

OpenStudy (raden):

you're welcome :)

OpenStudy (anonymous):

wait a minute so the 1 is not a prime number?

OpenStudy (anonymous):

wow! the 1 is not a prime number